OpenStudy (anonymous):

plz help me i will give medal Given the function f(x)=x/2+3 and g(x)3/4x-2, find the values of 2f(4)-3g(6).

OpenStudy (anonymous):

f (4)=4/2+3=5and g (6)=3/(4*6)-2 Then put the values in the equation

OpenStudy (anonymous):

ya thank u i got the answer